# WRF-GS: Wireless Radiation Field Reconstruction with 3D Gaussian Splatting

### v2

@@ IV-D Electromagnetic Splatting
In simple terms, we consider each projected Gaussian as a virtual TX that can affect multiple angles, while signals from multiple virtual TXs are received at each angle. Depending on the distance of these signals to the RX, they will be attenuated to different degrees, ultimately resulting in the received signal on each angle. As explained in Section III-A, the reconstructed spatial spectrum can be represented as a spatial power distribution function, where the value is proportional to the square of the amplitude of the received signal. Therefore, by deriving the received signal on each angle from Eqn. (12) and Eqn. (13), we obtain the spatial spectrum described in Eqn. (4), which serves as a representation of the underlying WRF.
IV-E Summary
The WRF-GS framework integrates physical and deep learning models to enable comprehensive environment-aware channel modeling and high-fidelity spatial spectrum synthesis. At the core of this approach lies the establishment of a connection between virtual TXs and 3D Gaussians. The scenario representation network models the spatial distribution of signal propagation and attenuation using 3D Gaussians based on input data. To transform these 3D spatial representations onto a 2D plane for further processing, we employ the Mercator map projection, ensuring accurate spatial mapping. Finally, the electromagnetic splatting module employs a parallel and efficient algorithm to compute pixel values in each direction, thereby achieving fast spatial spectrum synthesis.
V WRF-GS Implementation and Evaluation
@@
In this section, we implement and evaluate WRF-GS in a laboratory environment using a real-world dataset in [6].
V-A Implementation
WRF-GS is implemented in a Python environment using modified custom CUDA kernels for rasterization [10] and is trained on an NVIDIA GeForce RTX 3090 GPU. Detailed settings and implementation are provided below.
Position Initialization: We use the LiDAR 3D point clouds [^1] Note that WRF-GS also works well by randomly initializing the point clouds within a specified range of space. shown in Fig. 2(a) to initialize the positions of 3D Gaussian points. The reason is that the most significant signal attenuation usually occurs on the object’s surface [7], which corresponds to the 3D point clouds. This initialization enables a fast convergence of WRF-GS. Besides, we adopt an adaptive density control strategy [10] to learn the position distribution of the 3D Gaussians during the training process.
@@ V-A Implementation
where $\boldsymbol{t}$ represents the 3D coordinates of the TX and $L$ is the order of the position encoding. For the input position encoding of $P_{\text{TX}}$ and $G(\boldsymbol{x})$ in the MLPs, we set $L=9$.
Optimization Details: In the scenario representation network, we need to train two MLPs and learn the 3D Gaussian representations using the training dataset. We adopt the Adam optimizer for network training and 3D Gaussian representation. The loss function is defined as the difference between the synthesized spatial spectrum $I_{\text{pred}}$ and the ground truth $I_{\text{gt}}$, i.e.,
@@
where the $\xi(I_{\text{gt}},I_{\text{pred}})$ is the structural similarity index measure (SSIM) function [23], which is used to measure the similarity between two images. In addition, $\eta=0.2$ is a weighting factor.
V-B Experiment Design
Dataset. We use an open-source dataset provided in NeRF2 [6] to evaluate the WRF-GS framework. This dataset was obtained in a laboratory environment, which consists of multiple tables, shelves, and several small rooms. The 3D LiDAR point clouds of the environment are shown in Fig. 2(a). The RX is equipped with a $4\times 4$ antenna array and operated at 915MHz, while the TX is an RFID tag that sends RN16 messages repeatedly. To collect the training dataset, the position of the RX is fixed and the position of the TX is systematically varied throughout the laboratory. Each sample in the dataset consists of the TX position and the corresponding spatial spectrum measured at the RX (i.e., Eqn (4)).
Benchmarks. We compare WRF-GS with several baselines:
  • Ray Tracing[2]: This approach emits rays from the RX through every pixel in the laboratory using the 3D LiDAR point clouds. These rays interact with objects to simulate the propagation of signal in the real world.
@@ V-B Experiment Design
  • NeRF2 [6]: NeRF2 is a method for spatial spectrum synthesis using the NeRF technique. After training, NeRF2 can predict the spatial spectrum at arbitrary TX positions.
V-C Evaluation
@@
We first evaluate the accuracy of the spatial spectrum synthesis generated by the WRF-GS, NeRF2, VAE, and ray tracing methods, as shown in Fig. 2(c). The visualization is obtained by transforming the matrix form in Eqn. (4) into a polar coordinate representation. The spatial spectrum of WRF-GS for the case when the TX is located at position P3 is a projection of the constructed WRF visualized in Fig. 2(b). The results in Fig. 2(c) show that WRF-GS achieves the best prediction accuracy, closely matching the ground truth.
Fig. 10 compares the SSIM [23] values of the synthesized spatial spectra generated by different methods. SSIM values range from 0 to 1, with higher values indicating greater similarity between the synthesized and groud truth samples. The cumulative distribution function (CDF) of SSIM illustrates the disparities between the methods and the ground truth. From Fig. 10, we see that WRF-GS outperforms the baselines. Specifically, the median SSIM values for WRF-GS, NeRF2, VAE, DCGAN, and ray tracing are 0.82, 0.78, 0.70, 0.56, and 0.38, respectively. The slightly inferior performance of NeRF2 is due to the challenge in enumerating the virtual TX locations along the propagation paths for integration or rendering. In contrast, the ray tracing method achieves the worst performance, as it heavily relies on accurate environmental material information.
@@
Figure 10: The SSIM values of the synthesized spatial spectra generated by the WRF-GS, NeRF2, VAE, DCGAN, and ray tracing methods.
@@
Figure 11: The median SSIM values of the WRF-GS and NeRF2 methods under different training dataset sizes.
@@
Fig. 11 presents the median SSIM values of the WRF-GS and NeRF2 methods across varying sizes of the training dataset. We maintain a consistent test dataset and randomly subsample the original training dataset to create new training datasets ranging from $200$ to $4,800$ samples. The results demonstrate that the WRF-GS framework consistently outperforms the NeRF2 method for all training dataset sizes, improving the median SSIM by $4\%$ to $8\%$. More importantly, the performance gap increases as the training dataset size decreases. This suggests that WRF-GS is well-suited for practical applications in channel modeling, as it requires less training data and, consequently, less time and effort from experienced engineers.
@@
Figure 12: The rendering time required to synthesize a spatial spectrum using different methods under the same settings.
@@
Fig. 12 illustrates the rendering time required to synthesize a spatial spectrum using different methods under the same settings. The results show that, for processing a single sample, WRF-GS takes $0.005$s, while NeRF2, VAE, DCGAN, and Ray Tracing takes $0.2$s, $0.05$s, $0.09$s, and $7$s, respectively. The processing time of WRF-GS is usually one to three orders of magnitude less than other methods. This superior computational efficiency makes WRF-GS well-suited for latency-sensitive applications.
VI Case Study: CSI Prediction
@@
In this section, we demonstrate the effectiveness of WRF-GS on the downlink CSI prediction task in MIMO systems. Typically, the downlink and uplink transmission are operated on different frequency bands in frequency domain duplex systems, where the uplink-downlink reciprocity does not hold. The base station (BS) can obtain the downlink CSI of each antenna from the uplink channel feedback by sending pilot sequences to the end devices. However, the overhead of such feedback is linearly proportional to the number of antennas and devices, making it impossible in massive MIMO systems.
@@
Motivated by the fact that the uplink and downlink transmissions experience the same physical environment, it is reasonable to use the uplink CSI measured at the BS to infer the downlink CSI. To achieve this goal, we employ WRF-GS to predict the downlink CSI based on the uplink CSI. As pointed out in [27], each CSI is unique and highly correlated with the physical environment. Therefore, there is a mapping between the device’s location and its uplink CSI. Accounting for this, Eqn. (8) can be rewritten as
@@
where $I_{u}(\boldsymbol{x})$ is the uplink CSI at location $\boldsymbol{x}$. Thus, we can retrain a WRF-GS to perform the downlink CSI prediction task. In this case, the input is changed from the TX’s location to the uplink CSI, and the output is changed from the spatial spectrum to the downlink CSI.
VI-A Experiment Setup
We adopt the public Argos channel dataset [28] to train the MLPs and 3D Gaussian representations in WRF-GS for the downlink CSI prediction task. This dataset is collected in a real environment and the CSI is measured in different environments where a BS is equipped with $104$ antennas and serves multiple users. Each CSI measurement contains $52$ subcarriers. Similar to previous works [6, 17], we regard the first $26$ subcarriers as the uplink channel and the remaining $26$ subcarriers as the downlink channel. Apart from NeRF2 [6], we also compare WRF-GS with the following three approaches:

### v4

@@ V-A Scenario Representation Network Using Deformable 3D Gaussians
The dynamic components, which correspond to small-scale fading, are modeled using a deformation network ${D}_{\Theta}$. This network captures the variations in signal propagation paths caused by complex multipath effects, leading to changes in the coverage and shape of virtual TXs. The deformation network, implemented as an MLP, maps the coordinates of the 3D Gaussians and the TX position to their corresponding deviations in signal strength, rotation, and scaling. The weights $\Theta$ of the MLP are optimized through this mapping process. As shown in Fig. 11, the deformation network processes the input through eight fully connected layers with ReLU activations and 256-dimensional hidden layers, producing a 256-dimensional feature vector. This vector is then passed through three additional fully connected layers (without activation functions) to separately output the offsets for signal strength, rotation, and scaling. Notably, similar to NeRF, we concatenate the feature vector output from the fourth layer with the input and feed them into the fifth layer to enhance feature representation. Given the TX position, it is fed into the network along with the 3D Gaussians to generate the offsets:
${D}_{\Theta}:(G(\boldsymbol{x}),P_{\text{TX}})\Rightarrow\left(\Delta_{\text{sig}}(\boldsymbol{x}),\Delta_{\text{rot}}(\boldsymbol{x}),\Delta_{\text{scal}}(\boldsymbol{x})\right),$ (14)
where $\Delta_{\text{sig}}(\boldsymbol{x})$, $\Delta_{\text{rot}}(\boldsymbol{x})$, and $\Delta_{\text{scal}}(\boldsymbol{x})$ represent the offsets in signal strength, rotation, and scaling, respectively. Once the dynamic components are obtained, they are combined with the static components to form the complete properties of the 3D Gaussians at a specific TX position. These properties are then used in the subsequent coordinate projection transformation and electromagnetic signal calculation processes.
By incorporating deformable 3D Gaussians, the improved scenario representation network effectively integrates the characteristics of large-scale and small-scale fading into the channel model. This enhancement enables the network to accurately capture the high-frequency variations, significantly improving its interpretability and characterization capabilities.
Figure 10: The enhanced scenario representation network. The random 3D points are used to initialize 3D Gaussians, each characterized by properties such as position, opacity, signal strength, rotation, and scale. The 3D Gaussians, along with the TX position, are fed into the deformation network to compute dynamic offsets for the properties. These dynamic offsets are combined with the original static properties to derive the final characteristics of 3D Gaussians. These updated properties are then processed through the projection model and electromagnetic splatting to synthesize the final spatial spectra.
V-B Electromagnetic Splatting with $\alpha$-blending
With the introduction of a new structure in the scenario representation network, WRF-GS+ enhances the original framework by associating environmental information with the opacity attribute of 3D Gaussians. This opacity is translated into the weight of each virtual TX, which implicitly characterizes attenuation information, enabling a more accurate representation of signal propagation. Compared with WRF-GS, WRF-GS+ eliminates the need for additional attributes. This simplifies the 3D-GS model by reducing the number of parameters and memory overhead while improving convergence.
In the field of optics, opacity is modeled as particle density, which characterizes the probability of particles in space being blocked. This concept is adapted to the wireless domain, where opacity is used to represent the environmental influence on signal propagation. Specifically, the opacity of a 3D Gaussian reflects the material properties at its location, which in turn determines the weight of the corresponding virtual TX. By incorporating this property, WRF-GS+ can effectively model how environmental factors, such as obstacles or reflectors, affect signal strength at different locations.
Based on the $\alpha$-blending technique described in Eqn. (7) for optical rendering, we extend this concept to the electromagnetic domain by adapting it to complex number calculations. This allows us to rewrite Eqn. (12) and Eqn. (13) as:
$R_{k}=\displaystyle\sum_{i=1}^{N}({S}(\boldsymbol{x}_{i})+\Delta_{\text{sig}}(\boldsymbol{x}_{i})){\alpha}_{i}\displaystyle\prod_{j=1}^{i-1}(1-{\alpha}_{j}),$ (15)
where $R_{k}$ represents the received signal at a specific angle (pixel) $k$. This formula indicates that the received signal is the superposition of contributions from Gaussians, weighted by their respective $\alpha$-blending coefficients. Each Gaussian’s contribution consists of a static component ${S}(\boldsymbol{x}_{i})$ and a dynamic component $\Delta_{\text{Sig}}(\boldsymbol{x}_{i})$, reflecting the influence of environmental and TX movement effects, respectively.
To provide an intuitive understanding of this process, we revisit the example illustrated in Fig. 9(c). For the pixel value $R_{1}$ in the upper-left corner, only the signal from $G(\boldsymbol{x}_{1})$ contributes, but it is now weighted by its opacity coefficient $\alpha_{1}$. For the pixel value $R_{2}$ in the upper-right corner, the contributions from $G(\boldsymbol{x}_{1})$, $G(\boldsymbol{x}_{2})$, and $G(\boldsymbol{x}_{3})$ must be considered. The coefficient for the signal from $G(\boldsymbol{x}_{1})$ remains $\alpha_{1}$, while the coefficients for $G(\boldsymbol{x}_{2})$ and $G(\boldsymbol{x}_{3})$ are adjusted to $\alpha_{2}(1-\alpha_{1})$ and $\alpha_{3}(1-\alpha_{1})(1-\alpha_{2})$, respectively. The calculation for other pixel values $R_{k}$ follows the same principle.
By leveraging the static nature of environmental information, we associate this information with the opacity attribute of 3D Gaussians, which depends only on their positions. WRF-GS+ achieves a more accurate and interpretable representation of the WRF by decomposing the virtual TX into static and dynamic components.
Figure 11: The architecture of the deformation network, designed to learn the dynamic components in the scene. The 3D Gaussians and the TX position are first encoded using positional encoding and then fed into the deformation network. The network outputs the offsets in signal strength, rotation, and scaling attributes for each 3D Gaussian, enabling the representation of dynamic signal variations caused by TX movement and environmental interactions.
V-C Summary
The proposed WRF-GS+ framework addresses the limitations of WRF-GS in modeling dynamic signal variations and parameter redundancy. First, leveraging insights from the different characteristics of large-scale fading and small-scale fading in wireless communication scenarios, we enhance the scenario representation network by introducing deformable 3D Gaussians. This approach can efficiently capture the high-frequency signal variations caused by complex multipath effects. By integrating deep learning with physics-based modeling, the network achieves a more effective representation of high-frequency signal variations. Second, we enhance the electromagnetic splatting process by using the inherent opacity of 3D Gaussians as the weighting factor, eliminating the need for the network to explicitly represent attenuation parameters. This modification not only reduces computational complexity but also enhances the framework’s efficiency by leveraging the natural properties of 3D Gaussians to model signal attenuation implicitly. Finally, we implement an $\alpha$-blending method to compute the received signals from each direction in parallel, synthesizing the spatial spectrum with improved efficiency.
VI Implementation and Evaluation
@@
In this section, we implement and evaluate the WRF-GS and WRF-GS+ methods in a laboratory environment using a real-world dataset in [7].
VI-A Implementation Details
The proposed methods are implemented in Python, utilizing customized CUDA kernels for rasterization [11], and trained on an NVIDIA GeForce RTX 3090 GPU. It is worth noting that we handle complex signals by separating them into real and imaginary components via Euler’s formula rather than decomposing complex signals into amplitude and phase. This operation enables parallelized computations in the CUDA kernels. The implementation details are provided below.
Position Initialization: We randomly generate 3D points within a specified range. The positions of these points are used to initialize the center coordinates of the 3D Gaussians. In addition, we implement an adaptive density control strategy [11] to learn the positional distribution of the 3D Gaussians during the training process. Compared with the LiDAR point-cloud initialization [1], the random point initialization can overcome the requirement for multimodal input and minimize the reliance on the quality of LiDAR point clouds.
@@ VI-A Implementation Details
where $\boldsymbol{t}$ represents the 3D coordinates input and $L$ is the order of the position encoding. For the input position encoding of $P_{\text{TX}}$ and $G(\boldsymbol{x})$ in the MLPs, we set $L=9$.
Optimization Details: In the scenario representation network, we need to train the MLPs and learn the 3D Gaussian representations using the training dataset. We adopt the Adam optimizer for training and 3D Gaussian representation. The loss function is computed by comparing the difference between the synthesized spatial spectrum $I_{\text{pred}}$ and the ground truth $I_{\text{gt}}$:
@@
where $\xi(I_{\text{gt}},I_{\text{pred}})$ is the structural similarity index measure (SSIM) function [28], which is used to measure the similarity between two images. In addition, $\eta=0.2$ is a weighting factor.
VI-B Experiment Design
Dataset. We use an open-source dataset provided in NeRF2 [7] to evaluate the proposed frameworks. This dataset is obtained in a laboratory environment, which consists of multiple tables, shelves, and several small rooms. The 3D LiDAR point clouds of the environment are shown in Fig. 2(a). The RX is equipped with a $4\times 4$ antenna array and operated at 915MHz, while the TX is an RFID tag that sends RN16 messages repeatedly. To collect the training dataset, the position of the RX is fixed and the position of the TX is systematically varied throughout the laboratory. Each sample in the dataset consists of the TX position and the corresponding spatial spectrum measured at the RX (i.e., Eqn. (4)).
Benchmarks. We compare WRF-GS+ and WRF-GS with several baselines:
  • Ray Tracing[3]: This approach emits rays from the RX through every pixel in the laboratory using the 3D LiDAR point clouds. These rays interact with the objects, resulting in reflection, diffraction, and scattering, which are recursively traced to simulate real-world ray propagation using the RayTracing toolkit in MATLAB.
@@ VI-B Experiment Design
  • NeRF2 [7]: NeRF2 is a method for spatial spectrum synthesis using the NeRF technique. Similarly to our work, NeRF2 can synthesize the spatial spectrum at arbitrary TX positions after training in a given scenario.
VI-C Evaluation
@@
We first compare the accuracy of the spatial spectrum synthesis generated by the WRF-GS+, WRF-GS, NeRF2, VAE, and ray tracing methods, as shown in Fig. 2(b). The visualization is obtained by transforming the matrix form in Eqn. (4) into a polar coordinate representation. From Fig. 2(b) we see that WRF-GS+ achieves the best performance compared with the WRF-GS method and the baselines, closely matching the ground truth.
Next, we compare the SSIMs of the synthesized spatial spectra generated by different methods. The value of SSIM ranges from 00 to $1$, where a higher value indicates greater similarity between the synthesized and ground truth samples. The cumulative distribution function (CDF) of SSIM illustrates the disparities between the methods and the ground truth. From Fig. 12, we see that the median SSIM values for WRF-GS+, WRF-GS, NeRF2, VAE, DCGAN, and ray tracing are $0.90$, $0.82$, $0.78$, $0.70$, $0.56$, and $0.38$, respectively, and their 90th percentiles are $0.95$, $0.88$, $0.86$, $0.82$, $0.69$, and $0.61$. Evidently, WRF-GS+ significantly outperforms WRF-GS thanks to its adaptation of deformable 3D Gaussians in the scenario representation network design. This operation enables the WRF-GS+ method to efficiently characterize the distribution of the WRF in a given scene by separately modeling the static and dynamic components of the signals from virtual TXs. In addition, the poor performance of NeRF2 is due to its reliance on the voxel sampling rate along the pixel ray. As a result, a high resolution results in significant computational overhead and demands extensive training data, while a lower resolution compromises the representation of WRF. Thus, there is a trade-off between representation capability and computational efficiency. In contrast, the ray tracing method achieves the poorest performance as it heavily relies on environmental material information, which is often hard to obtain.
@@
Figure 12: The SSIM values of the synthesized spatial spectra generated by the WRF-GS+, WRF-GS, NeRF2, VAE, DCGAN, and ray tracing methods.
@@
Figure 13: The median SSIM values of the WRF-GS+, WRF-GS and NeRF2 methods under different training dataset sizes.
@@
Fig. 13 compares the median SSIM values of the WRF-GS+, WRF-GS, and NeRF2 methods across varying sizes of the training dataset. We maintain a consistent test dataset and randomly subsample the original training dataset to create new training datasets ranging from $200$ to $4,800$ samples. It can be seen that WRF-GS+ consistently outperforms the WRF-GS and NeRF2 methods for all training dataset sizes, improving the median SSIM by $3\%$ to $13\%$ and $11\%$ to $19\%$, respectively. This demonstrates that WRF-GS+ has the highest sample efficiency. More importantly, the performance gap increases as the training dataset size decreases, which indicates that the proposed methods require less training data and, consequently, less time and effort from experienced engineers.
@@
Figure 14: The rendering time required to synthesize a spatial spectrum using different methods.
@@
Fig. 14 illustrates the rendering time required to synthesize a spatial spectrum using different methods under the same settings. The results show that WRF-GS+ achieves a processing time of $0.008$s per sample, while WRF-GS, NeRF2, VAE, DCGAN, and Ray Tracing require $0.005$s, $0.2$s, $0.05$s, $0.09$s, and $7$s, respectively. Notably, the processing times of WRF-GS+ and WRF-GS are one to three orders of magnitude lower than those of the other methods. This significant advantage makes the proposed methods highly suitable for latency-sensitive applications. Although the processing time of WRF-GS+ is slightly higher than that of WRF-GS, this difference is primarily due to the inclusion of the deformation network in WRF-GS+, which introduces more parameters compared to the two MLPs used in WRF-GS. However, given the significant improvement in spectrum synthesis accuracy achieved by WRF-GS+, this millisecond-level difference is negligible and acceptable in most practical application scenarios.
VII Case Studies
In this section, we further demonstrate the effectiveness of the proposed methods on RSSI and CSI prediction tasks.
VII-A Case Study I: RSSI Prediction
We first apply the WRF-GS and WRF-GS+ methods to the RSSI prediction task, which involves estimating the signal strength at a specific location based on factors such as distance, environmental conditions, and prior measurements. RSSI prediction is essential for network management, indoor positioning, and signal coverage planning. For example, the RSSI prediction accuracy directly affects the positioning accuracy of the localization task and the performance of the base station (BS) deployment in cellular networks.
Unlike the previous spatial spectrum reconstruction task, the RSSI measurements are obtained by a single omnidirectional antenna instead of an antenna array. As a result, the output is a composed signal containing the power from all directions rather than the $90\times 360$ pixel matrix. To complete the RSSI prediction task, the proposed methods are trained on the public BLE dataset [7]. We compare the proposed methods with the NeRF2 method and the MRI [19] method. MRI is a learning-based method that interpolates the RSSI values at the unsampled location using a log-distance path-loss model.
VII-A1 Experiment Setup
The BLE dataset is obtained in a scene containing several rooms with a total area of about $15,000ft^{2}$. There are $21$ RXs and a TX moving in this area. The RSSI at these RXs is recorded when the TX moves from one position to another. There are a total of $6,000$ RSSI measurements in the BLE dataset. Note that we consider the measurement to be an invalid value when the RX is far from the TX. In this situation, the recorded RSSI is set to $-100$ dBm. There are $2,000$ valid RSSI measurements for each RX, where $1,600$ measurements are used for training and the remaining $400$ are used for testing.
VII-A2 RSSI Prediction Accuracy
Figure 15: The RSSI prediction errors of the WRF-GS+, WRF-GS, NeRF2, and MRI methods on the public BLE dataset.
Fig. 15 compares the RSSI prediction errors of the WRF-GS+, WRF-GS, NeRF2, and MRI methods on the public BLE dataset. The prediction error is defined as the absolute difference between the predicted and ground truth RSSI values. In simulation, we train the proposed methods on the dataset of each RX. The result measured in dB is the average value of the RSSI prediction errors for $12$ RXs. From Fig. 15, we see that WRF-GS+ has the lowest RSSI prediction error among the WRF-GS, NeRF2, and MRI methods. The median of WRF-GS+ is $2.4$ dB ($10$-th percentile: $1.1$ dB; $90$-th percentile: $4.6$ dB). In contrast, the median errors of WRF-GS, NeRF2, and MRI are $2.9$ dB, $3.1$ dB, and $8.3$ dB, respectively. The proposed methods clearly outperform the baselines, indicating that the 3D-GS-based method is more suitable than the NeRF-based method in this task. The superior performance of the WRF-GS+ is due to its ability to separate the signals from virtual TXs into static and dynamic components. This operation enables an accurate RSSI prediction even with continuous TX movements.
VII-B Case Study II: CSI Prediction
@@
We adapt the WRF-GS+ and WRF-GS methods to the downlink CSI prediction task in MIMO systems. Typically, the downlink and uplink transmission are operated on different frequency bands in frequency domain duplex systems, where the uplink-downlink reciprocity does not hold. The BS can obtain the downlink CSI of each antenna from the uplink channel feedback by sending pilot sequences to the end devices. However, the overhead of such feedback is linearly proportional to the number of antennas and devices, making it impossible in massive MIMO systems.
@@
Motivated by the fact that the uplink and downlink transmissions experience the same physical environment, it is reasonable to use the uplink CSI measured at the BS to infer the downlink CSI. To achieve this goal, we employ WRF-GS to predict the downlink CSI based on the uplink CSI. As pointed out in [32], each CSI is unique and highly correlated with the physical environment. Therefore, there is a mapping between the device’s location and its uplink CSI, which is similar to fingerprint-based localization. Accounting for this, Eqn. (8) can be rewritten as
@@
where $I_{u}(\boldsymbol{x})$ is the uplink CSI at location $\boldsymbol{x}$. Thus, we can apply the WRF-GS framework to the downlink CSI prediction task by employing a similar configuration in Section VI-A.
For the WRF-GS+ method, we adapt the deformation network to this CSI prediction task by modifying Eqn. (14) into
${D}_{\Theta}:(G(\boldsymbol{x}),I_{u}(\boldsymbol{x}))\Rightarrow\left(\delta_{\text{Sig}}(\boldsymbol{x}),\delta_{\text{r}}(\boldsymbol{x}),\delta_{\text{s}}(\boldsymbol{x})\right).$ (19)
Note that the input is changed from the TX’s location to the uplink CSI, and the output of the whole pipeline is changed from the spatial spectrum to the downlink CSI.
Figure 16: Comparison of the ground truth CSI and the predicted CSI using the proposed methods in terms of amplitude and phase via different sub-carriers.
VII-B1 Experiment Setup
We train the WRF-GS and WRF-GS+ methods on the public Argos channel dataset [33] for the downlink CSI prediction task. This dataset is collected in a real environment. The CSI is measured in different environments where a BS is equipped with $104$ antennas and serves multiple users. Each CSI measurement contains $52$ subcarriers. Similarly to [7] and [20], we regard the first $26$ subcarriers as the uplink channel and the remaining $26$ subcarriers as the downlink channel. Unlike the spatial spectrum reconstruction task, which requires computing power from all directions separately, we focus on a complex value of the predicted CSI at each subcarrier. Notice that this complex value is a composed signal that sums in all directions. Apart from NeRF2, we also compare the proposed methods with the following baselines: